OpenStudy (anonymous):

Solve the system of equations. 3x - 4y = -10 and 5x + 8y = -2

OpenStudy (anonymous):

\[3x-4y=-10\]\[5x+8y=-2\] Multiply by 2 the first equation and sum with the second: \[2(3x-4y)=2(-10)\] 6x-8y=-20....+ 5x+8y=-2 ------------- 11x=-22 x=-1 replace x in some equation: 3(-1)-4y=-10 -3-4y=-10 y=7/4

OpenStudy (anonymous):

(1, 2) (2, 1) (1, -2) (-2, 1)

OpenStudy (anonymous):

its a multiple answer question

OpenStudy (anonymous):

i want you to understand the idea said ... the first equation is Perfectly the same as second in the points of intersection and we have to equation just to know variable x and y . Okey ? if we draw first equation it would like so similar to this :  the second equation :  so you have points and it may make either equation 1 or 2 correct .. its clear that or maybe both of them .. in your Q ,, both of them correct : the choice : ( -2,1)

OpenStudy (anonymous):

sorry i have to mistake in English : second line to = two .. and at last statement ... ITs clear that maybe both of them .

OpenStudy (anonymous):

no worries my friend. your english is just fine! :)

OpenStudy (anonymous):

aha i read it, now i have a better picture of how to do it